OpenStudy (anonymous):

3z+4z+b-5? Medals included...Always:)

jigglypuff314 (jigglypuff314):

-.- terms with same variables can add/subtract together what is 3 + 4 = ? then what is 3z + 4z = ?

jigglypuff314 (jigglypuff314):

because I've seen that you've posted similar questions, only to get direct answers (which is against the rules) and that you still haven't actually learned how to do it...

jigglypuff314 (jigglypuff314):

I would suggest not giving medals to those who give direct answers because it is against the rules and if a moderator finds out, they can report to your teachers/school by giving such medals would be encouraging cheating.

OpenStudy (superdavesuper):

Sorry jigglypuff I am new here and did not see the code of conduct before. I will not give out the answers again.

jigglypuff314 (jigglypuff314):

I'm glad you understand now :D

jigglypuff314 (jigglypuff314):

for questions like yours @MariLearningHW, it's the concept of apples and oranges (as my teachers like to say) you can't compare apples and oranges, they're two different things. but you can add apples with apples :) so same with variables. a + 2a = 3a

OpenStudy (anonymous):

I just joined today...I didnt know either

jigglypuff314 (jigglypuff314):

it's alright :) when I first joined I hadn't know either, it's all in the learning experiance :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Can a ask a math question now?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Charlie has x state quarters. Ty has 3 more quarters than Charlie has. Vinnie has 2 times as many quarters as Ty has. Write and simplify on expression to show how many state quarters they will have in all.

jigglypuff314 (jigglypuff314):

"Charlie has x state quarters" so C = x "Ty has 3 more quarters than Charlie has" has means equals, more means adding so T = 3 + C "Vinnie has 2 times as many quarters as Ty" "times" means multiplying so V = 2 T how many they would have in all would be C + T + V follow so far? :)